Convert 9.52 x 104 g O to atoms of Ca(OH)2


1
Expert's answer
2021-02-02T01:22:09-0500

Molar mass of O is 16 gmol-1

16 g atom means 1 mol O atoms = 6.023 × 1023 no of O atoms.

So 9.52 × 104 g O means = (9.52 × 104)×(6.023× 1023) \ 16 = 3.58 × 1027 atoms of O.

Here Ca(OH)2 is a compund so it is present as unit of calcium hydroxide not as atoms

So in one Ca(OH)2 2 O atoms present

That means 2 O atoms present in 1 Ca(OH)2

So 3.58 × 1027 O atoms present in

= 3.58 × 1027 \ 2 = 1.79 × 1027 Ca(OH)2 unit.

That means Ans is 1.79× 1027 .